Peter walks 10 miles in 3 hours. How long would it take him to walk 1 mile ?
Answer by lenny460(1073)

Peter walks 10 miles in 3 hours. How long would it take him to walk 1 mile ?
1 hour = 60 minutes
3 hours = 60 * 3 = 180 minutes
Therefore:
Peter walks 10 miles in 180 minutes
10 = 180
Divide each side by 10
10/10 = 180/10
1 = 18

Answer:
Peter will take 18 minutes to walk 1 mile.










Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: lennoxobuong@yahoo.com